OpenStudy (anonymous):

when 8^263 is expanded out,what digit will be in the units (1's) space? a)8 b)4 c)2 d)6 e)5 ???

OpenStudy (anonymous):

3255866422304616344765552632188114158762089024568314531443485259650408807528140659922574316831813618526821245406949824436469141432675471230646169947427163108444901161872077421124549944292877941762189949285100879873872435565174053364826112 It is a two

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

there are a couple of ways to do this. you can find the pattern in the powers. 8^1 gives a ones digit of 8 8^2 gives a ones digit of 4 8^3 gives 2 8^4 gives 6 8^5 gives 8 and it will repeat from there. So we have a sequence like: 8, 4, 2, 6, 8, 4, 2, 6, .... and we want to know what the 263rd term is. Since this cycles every four terms, we do 263/4 and find its remainder. \[263 = 4(65)+3\]the remainer is 3, so the 263rd term will land on the third term of the cycle which is 2.

OpenStudy (anonymous):

joemath will u help me in a problem?

OpenStudy (anonymous):

i can try, whats up?

OpenStudy (anonymous):

0/0=undefined 2*0=0 2=0/0 then,2=undefined? but it is impossible

OpenStudy (anonymous):

right, thats impossible. because 0/0 is undefined, it wont follow any of our rules we have in math. Its like all those proofs you see for 1 = 0 and the like.

OpenStudy (anonymous):

if that were a true statement, then not just 2, but any number would be undefined. \[a*0 = 0 \Rightarrow a = \frac{0}{0}\]

OpenStudy (anonymous):

yes not only 2.so,it does not follow the rules.

OpenStudy (anonymous):

right. whenever someone comes to you with some bogus claim like: \[\sqrt{-4} = 2\] or whatever (that happened yesterday lol), they are always breaking a rule somewhere, the name of the game is, can you find what they did wrong.

OpenStudy (anonymous):

ok.thanks

OpenStudy (anonymous):

The only sorta "weird" statement that is a true statement is: \[.9999\ldots = 1\] That is true. Other than that though, all bogus! lolol

OpenStudy (anonymous):

maths is full of mystery and beauty .....

OpenStudy (anonymous):

indeed it is....indeed it is lol.

OpenStudy (anonymous):

lol